# Green currents of holomorphic correspondences on compact Kähler manifolds

### 06 Mar 2026 v1

@@ 1. Introduction
Under some mild assumptions on the local multiplicities of the graph of a holomorphic correspondence, (an analogue of not having highly critical periodic orbits in dimension $1$), our second main result strengthens the results of [19]. See also [37] for the case of dimension 1 and [28, Corollary 4.8] for polynomial correspondences. More precisely, we show that the convergence towards Green currents is exponentially fast. In the case of automorphisms of Kähler manifolds, this property has been the key to establish many strong statistical properties of the dynamical system, see e.g. [11, 25, 45]. We also show that these assumptions are satisfied by generic holomorphic correspondences. In particular, this is true for polynomial correspondences up to finite iterations (see Corollaries 5.2 & 5.2). The definitions of $q$-small inverse multiplicity and small multiplicity are given in Definition 4.
@@
    If $f$ has $q$-small inverse multiplicity, then $d_{q}^{-n}(f^{n})^{*}(S)$ converges to $T^{+}$ exponentially fast;
@@
Acknowledgements. The first author has received funding from the National University of Singapore and MOE of Singapore through the grant A-8002488-00-00. The second author is part of the PHC Galileo project G24-123. A part of this work was done during the first author’s visit to the Centro di Ricerca Matematica Ennio De Giorgi and the University of Pisa. He would like to thank them for their hospitality and the excellent working conditions.
@@ 3.2. Regularity of Green currents
Lemma 3.3. Let $X_{1},X_{2}$ be compact Kähler manifolds, and let $A\subset X_{1}\times X_{2}$ be an analytic subset. Suppose that the canonical projection $\pi_{1}:X_{1}\times X_{2}\longrightarrow X_{1}$, when restricted to $A$, is a ramified covering of maximal multiplicity $\rho$ and with finite fibers of cardinality $d$ (counted with multiplicity).
There exists a constant $C=C(X_{1},X_{2},A)$ such that, for every $z\in X_{1}$, there is an open set $U\ni z$ of $X_{1}$ such that for every $x\in A$ with $\pi_{1}(x)\in U$ we have
@@
Moreover, for every $y,y^{\prime}\in U$ we can write
@@
It then suffices, for every $z^{\prime}\in X_{1}$, to choose $U_{z^{\prime}}=\displaystyle\bigcap_{w\in\pi_{1}^{-1}(z^{\prime})\cap A}\tilde{D}_{w}$. Observe that, since $\pi_{1}|_{V_{w}\cap A}$ is a ramified covering of degree $\rho_{w}$ and $\sum_{w\in\pi_{1}^{-1}(z^{\prime})\cap A}\rho_{w}=d$, the set $\bigcup_{w\in\pi_{1}^{-1}(z^{\prime})\cap A}\pi_{1}|_{V_{w}}^{-1}(z)\cap A$ contains exactly $d$ point (counted with multiplicity) for every $z\in U_{z^{\prime}}$, i.e., it coincides with $\pi_{1}^{-1}(z)\cap A$. It follows that (7) holds for every $z\in U_{z^{\prime}}$ and (8) holds for every $y,y^{\prime}\in U_{z^{\prime}}$, for constants $\tilde{C}=\displaystyle\max_{w\in\pi_{1}^{-1}(z^{\prime})\cap A}\{C_{w}\}$ and $\tilde{\rho}=\displaystyle\max_{w\in\pi_{1}^{-1}(z^{\prime})\cap A}\{\rho_{w}\}$ (we can always choose the neighbourhoods sufficiently small, so that $\mathrm{dist}(y,y^{\prime})<1$).
Since by compactness we just need to consider a finite number of $z^{\prime}$, and every $z^{\prime}$ has a finite number of points $w$ in its fiber over $A$, we can take as the constant $C$ the maximum of the constants $C_{w}$. Moreover, $\tilde{\rho}$ is less than or equal to the maximal multiplicity $\rho$ of $\pi_{1}$ when restricted to $A$. The proof is complete. ∎
@@
Definition 3.5. The local multiplicity $\rho:=\rho(f)$ of $f$ is the maximal multiplicity of $\pi_{1}|_{\Gamma}$ seen as a ramified covering. We also set $\kappa:=1/\big(25k(4\rho)^{q}\big)$.
@@
  • the image $f(U)$ is the union of small open sets, and there is a smooth function $\tilde{u}$ such that $\omega=\mathrm{dd^{c}}\tilde{u}$ in $f(U)$;
  • Lemma 3.2 applies in $U$ for $X_{1}=X_{2}=X$ and $A=\Gamma$.
@@
4. Inverse multiplicity and equidistribution results
@@
Definition 4.2. The inverse multiplicity of $f$ denoted by $\delta(f)$ is the maximal local multiplicity of $\pi_{2}|_{\Gamma}$. For $0<q\leq k$, we say that $f$ has $q$-small inverse multiplicity if $d_{q}(f)>d_{q-1}(f)$ and $\tilde{\kappa}(f)^{-1}:=25k(4\delta(f))^{k-q+1}<d_{q}(f)/d_{q-1}(f)$.
Since the graph of $f^{-1}$ is obtained from $\Gamma$ by involution, we have $\delta(f)=\rho(f^{-1})$.
Lemma 4.3. Let $f$ be a holomorphic correspondence on $X$ such that $f^{-1}$ is also holomorphic. Let $R$ be a smooth current in $\mathcal{D}_{k-q+1}^{0}$. Then $\mathcal{U}_{(f^{n})_{*}(R)}$ is $(C_{R},\tilde{\kappa}^{n})$-Hölder continuous for some constant $C_{R}>0$ independent of $n$.
@@
Fix a smooth $\alpha$-normalized potential $U_{R}$ of $R$. Take a current $\Omega\in\mathcal{D}_{q}^{0}$ with $\|\Omega\|_{*}\leq 1$. Since $R$ is smooth, $(f^{n})_{*}(R)$ has a continuous super-potential. So, by Lemma 4 applied to $f^{n}$ instead of $f$, we have
$\mathcal{U}_{(f^{n})_{*}(R)}(\Omega)=\mathcal{U}_{\Omega}((f^{n})_{*}(R))=\mathcal{U}_{(f^{n})^{*}(\Omega)}(R)=\langle(f^{n})^{*}(\Omega),U_{R}\rangle.$
Applying inductively Corollary 3.2 to $f^{-1}$, we obtain $\|(f^{n})^{*}(\Omega)\|_{-2}\leq C\|\Omega\|_{-2}^{\tilde{\kappa}^{n}}$ with $C=\max\{C_{1},1\}^{\frac{1}{1-\tilde{\kappa}}}$. Hence $|\mathcal{U}_{(f^{n})_{*}(R)}(\Omega)|\leq C\|U_{R}\|_{\mathcal{C}^{2}}\|\Omega\|_{-2}^{\tilde{\kappa}^{n}}$. This completes the proof. ∎
Proposition 4.4. Let $f$ be a holomorphic correspondence on $X$ with $q$-small inverse multiplicity. Let $S$ be a current in $\mathcal{D}_{q}^{0}$ with $\|S\|_{*}\leq 1$. Let $R$ be a smooth (kq,kq)(k-q,k-q)-current with $\|\mathrm{dd^{c}}R\|_{*}\leq 1$. Then
$|\langle d_{q}^{-n}(f^{n})^{*}(S),R\rangle|\leq C_{6}\lambda_{0}^{n}$
where $C_{6}>0$ is a constant independent of $S$ and $n$, and $0<\lambda_{0}<1$ can be any constant strictly larger than $d_{q-1}(d_{q}\tilde{\kappa})^{-1}$.
@@
Since $f$ has $q$-small inverse multiplicity, we can choose constants $d_{q-1}<\lambda^{\prime}<\lambda^{\prime\prime}<d_{q}$ such that $1/\tilde{\kappa}<d_{q}/\lambda^{\prime\prime}<d_{q}/\lambda^{\prime}<d_{q}/d_{q-1}$. By Lemma 4, we have
@@
When $n$ is large enough, we have $\|(f^{n})_{*}(\mathrm{dd^{c}}R)\|_{*}\lesssim(\lambda^{\prime})^{n}$. By Lemma 4, $\mathcal{U}_{(\lambda^{\prime})^{-n}(f^{n})_{*}(\mathrm{dd^{c}}R)}$ is $(C_{R}(\lambda^{\prime})^{-n},\tilde{\kappa}^{n})$-Hölder continuous. By Proposition 2.1, we have
$|\langle d_{q}^{-n}(f^{n})^{*}(S),R\rangle|\lesssim Ad_{q}^{-n}(\lambda^{\prime})^{n}(1+\tilde{\kappa}^{-n}\log^{+}(C_{R}(\lambda^{\prime})^{-n}))\lesssim(\lambda^{\prime\prime}\tilde{\kappa}^{-1}d_{q}^{-1})^{n}.$
@@
Corollary 4.6. Let $f$ be as in Proposition 4 and $S$ be a current in $\mathcal{D}_{q}$. If the sequence $n_{i}^{1-m}d_{q}^{-n_{i}}(f^{n_{i}})^{*}\{S\}$ converges to some class $c$, then $n_{i}^{1-m}d_{q}^{-n_{i}}(f^{n_{i}})^{*}(S)$ converges SP-uniformly to $T_{c}$.
@@
In the following, we assume $f$ is a holomorphic correspondence on $X$ with simple action on cohomology and $f^{-1}$ is also holomorphic. Suppose $d:=d_{q}(f)$ is the main dynamical degree of $f$. Since we have $(f^{-1})^{*}=f_{*}$ and by duality, we get $d_{s}(f^{-1})=d_{k-s}(f)$ for all $0\leq s\leq k$. In particular, $f^{-1}$ also has simple action on cohomology with the same main dynamical degree. Let $T^{+}$ (resp. $T^{-}$) be the unique Green current of order $q$ (resp. kqk-q) associated with $f$ (resp. $f^{-1}$) of mass 1 such that $f^{*}(T^{+})=dT^{+}$ (resp. $f_{*}(T^{-})=dT^{-}$). Note that the correspondences in the family considered in Corollary 5.2 all have simple action on cohomology. The following gives Part (1) of Theorem 1.
Corollary 4.7. Let $f$ be a holomorphic correspondence on $X$ with $q$-small inverse multiplicity and simple action on cohomology. Let $S$ be a current in $\mathcal{D}_{q}$ with $\langle T^{-},S\rangle=1$. Then for any smooth test (kq,kq)(k-q,k-q)-form $\varphi$, we have
$|\langle d^{-n}(f^{n})^{*}(S)-T^{+},\varphi\rangle|\leq C_{7}\|\varphi\|_{\mathcal{C}^{2}}\lambda_{1}^{n}$
for some constants $C_{7}>0$ and $0<\lambda_{1}<1$ that could depend on $S$, $_{q-1}, dd$ and $\tilde{\kappa}$, but are independent of $n$.
@@
Let $\alpha$ and $S^{\prime}$ be as in the proof of Corollary 4. Since the action of $f^{*}$ is simple on cohomology, by [1, Lemma 4.1], $d^{-n}(f^{n})^{*}(\alpha)$ converges exponentially fast to the Green current. The assertions follow from Proposition 4 applied with $S^{\prime}$ instead of $S$. ∎
Before we proceed, we briefly discuss the speed of the above convergence. In Proposition 4, we can take $\lambda_{0}$ to be any constant strictly larger than $d_{q-1}(d\tilde{\kappa})^{-1}$. Hence, by [1, Lemma 4.1] (or our proof of Proposition 3.1), $\lambda_{1}$ is determined by the gap between $d$ and the remaining eigenvalues of $f^{*}:H^{q,q}(X,\mathbb{R})\to H^{q,q}(X,\mathbb{R})$, $l_{q}=d/d_{q-1}$ and $d_{q-1}(d\tilde{\kappa})^{-1}$.
@@
Definition 4.10. Let $f$ be a holomorphic correspondence on $X$ such that $f^{-1}$ is also a holomorphic correspondence. We say that $f$ has small multiplicity if the holomorphic correspondence $F=(f,f^{-1})$ has $k$-small inverse multiplicity.

### 25 Mar 2026 v2

@@ 1. Introduction
Under some mild assumptions on the local multiplicities of the graph of a holomorphic correspondence, (an analogue of not having highly critical periodic orbits in dimension $1$), our second main result strengthens the results of [19]. See also [37] for the case of dimension 1 and [28, Corollary 4.8] for polynomial correspondences. More precisely, we show that the convergence towards Green currents is exponentially fast when tested against smooth forms. In the case of automorphisms of Kähler manifolds, this property has been the key to establish many strong statistical properties of the dynamical system, see e.g. [11, 25, 45]. We also show that these assumptions are satisfied by generic holomorphic correspondences. In particular, this is true for polynomial correspondences up to finite iterations (see Corollaries 5.2 & 5.2). The definitions of $q$-small adjoint multiplicity and small multiplicity are given in Definition 4.
@@
    If $f$ has $q$-small adjoint multiplicity, then $d_{q}^{-n}(f^{n})^{*}(S)$ converges to $T^{+}$ exponentially fast;
@@
Acknowledgements. The first author has received funding from the National University of Singapore and MOE of Singapore through the grant A-8002488-00-00. A part of this work was done during the first author’s visit to the Centro di Ricerca Matematica Ennio De Giorgi and the University of Pisa. He would like to thank them for their hospitality and the excellent working conditions.
@@ 3.2. Regularity of Green currents
Lemma 3.3. Let $X_{1},X_{2}$ be Kähler manifolds with $X_{1}$ compact, and let $A\subset X_{1}\times X_{2}$ be an analytic subset. Suppose that the canonical projection $\pi_{1}:X_{1}\times X_{2}\longrightarrow X_{1}$, when restricted to $A$, is a ramified covering of maximal multiplicity $\rho$ and with finite fibres of cardinality $d$ (counted with multiplicity).
There exists a constant $C=C(X_{1},X_{2},A)$ such that for every $z\in X_{1}$ we have
@@
Moreover, for every $y,y^{\prime}\in X_{1}$ we can write
@@
It then suffices, for every $z^{\prime}\in X_{1}$, to choose $U_{z^{\prime}}=\displaystyle\bigcap_{w\in\pi_{1}^{-1}(z^{\prime})\cap A}\tilde{D}_{w}$. Observe that, since $\pi_{1}|_{V_{w}\cap A}$ is a ramified covering of degree $\rho_{w}$ and $\sum_{w\in\pi_{1}^{-1}(z^{\prime})\cap A}\rho_{w}=d$, the set $\bigcup_{w\in\pi_{1}^{-1}(z^{\prime})\cap A}\pi_{1}|_{V_{w}}^{-1}(z)\cap A$ contains exactly $d$ point (counted with multiplicity) for every $z\in U_{z^{\prime}}$, i.e., it coincides with $\pi_{1}^{-1}(z)\cap A$. It follows that (7) holds for every $z\in U_{z^{\prime}}$ and (8) holds for every $y,y^{\prime}\in U_{z^{\prime}}$, for constants $\tilde{C}=\displaystyle\max_{w\in\pi_{1}^{-1}(z^{\prime})\cap A}\{C_{w}\}$ and $\tilde{\rho}=\displaystyle\max_{w\in\pi_{1}^{-1}(z^{\prime})\cap A}\{\rho_{w}\}$ (we can always choose the neighbourhoods sufficiently small, so that $\mathrm{dist}(y,y^{\prime})<1$ for every $y,y^{\prime}\in U_{z^{\prime}}$).
By compactness, we just need to consider a finite number of $z^{\prime}$, and every $z^{\prime}$ has a finite number of points $w$ in its fibre over $A$. Therefore, for (7) it suffices to take as $C$ the maximum of the constants $C_{w}$, and for (8) we take a sufficiently large constant to also deal with couples of points that are far apart, i.e., that do not belong to the same $U_{z^{\prime}}$. Moreover, $\tilde{\rho}$ is less than or equal to the maximal multiplicity $\rho$ of $\pi_{1}$ when restricted to $A$. The proof is complete. ∎
@@
Definition 3.5. The local multiplicity $\rho:=\rho(f)$ of $f$ is the maximal multiplicity of $\pi_{1}|_{\Gamma}$ seen as a ramified covering. Set $\kappa:=1/\big(25k(4\rho)^{q}\big)$.
@@
  • the image $f(U)$ is the union of small open sets, and there is a smooth function $\tilde{u}$ such that $\omega=\mathrm{dd^{c}}\tilde{u}$ in $f(U)$.
@@
4. Adjoint multiplicity and equidistribution results
@@
Definition 4.2. The adjoint multiplicity of $f$ denoted by $\delta(f)$ is the maximal local multiplicity of $\pi_{2}|_{\Gamma}$. For $0<q\leq k$, we say that $f$ has $q$-small adjoint multiplicity if $d_{q}(f)>d_{q-1}(f)$ and $\tilde{\kappa}(f)^{-1}:=25k(4\delta(f))^{k-q+1}<d_{q}(f)/d_{q-1}(f)$.
Since the graph of $f^{-1}$ is obtained from $\Gamma$ by involution $(x,y)\mapsto(y,x)$, we have $\delta(f)=\rho(f^{-1})$.
Lemma 4.3. Let $f$ be a holomorphic correspondence on $X$ such that $f^{-1}$ is also holomorphic. Let $R$ be a smooth (kq,kq)(k-q,k-q)-form with $\|R\|_{\mathcal{C}^{2}}\leq 1$. Then $\mathcal{U}_{(f^{n})_{*}(\mathrm{dd^{c}}R)}$ is $(C_{6},\tilde{\kappa}^{n})$-Hölder continuous for some constant $C_{6}>0$ independent of $R$ and $n$.
@@
Take a current $\Omega\in\mathcal{D}_{q}^{0}$ with $\|\Omega\|_{*}\leq 1$. Since $R$ is smooth, $(f^{n})_{*}(\mathrm{dd^{c}}R)$ has a continuous super-potential. So, by Lemma 4 applied to $f^{n}$ instead of $f$, we have
$\mathcal{U}_{(f^{n})_{*}(\mathrm{dd^{c}}R)}(\Omega)=\mathcal{U}_{\Omega}((f^{n})_{*}(\mathrm{dd^{c}}R))=\mathcal{U}_{(f^{n})^{*}(\Omega)}(\mathrm{dd^{c}}R)=\langle(f^{n})^{*}(\Omega),R\rangle.$
Applying inductively Corollary 3.2 to $f^{-1}$, we obtain $\|(f^{n})^{*}(\Omega)\|_{-2}\leq C_{6}\|\Omega\|_{-2}^{\tilde{\kappa}^{n}}$ for some $C_{6}>0$ independent of $\Omega$. Hence $|\mathcal{U}_{(f^{n})_{*}(\mathrm{dd^{c}}R)}(\Omega)|\leq C_{6}\|\Omega\|_{-2}^{\tilde{\kappa}^{n}}$. This completes the proof. ∎
Proposition 4.4. Let $f$ be a holomorphic correspondence on $X$ with $q$-small adjoint multiplicity. Let $S$ be a current in $\mathcal{D}_{q}^{0}$ with $\|S\|_{*}\leq 1$. Let $R$ be a smooth (kq,kq)(k-q,k-q)-form with $\|R\|_{\mathcal{C}^{2}}\leq 1$. Then
$|\langle d_{q}^{-n}(f^{n})^{*}(S),R\rangle|\leq C_{7}\lambda_{0}^{n}$
where $C_{7}>0$ is a constant depending only on $f$, and $0<\lambda_{0}<1$ can be any constant strictly larger than $d_{q-1}(d_{q}\tilde{\kappa})^{-1}$.
@@
Since $f$ has $q$-small adjoint multiplicity, we can choose constants $d_{q-1}<\lambda^{\prime}<\lambda^{\prime\prime}<d_{q}$ such that $1/\tilde{\kappa}<d_{q}/\lambda^{\prime\prime}<d_{q}/\lambda^{\prime}<d_{q}/d_{q-1}$. By Lemma 4, we have
@@
When $n$ is large enough, we have $\|(f^{n})_{*}(\mathrm{dd^{c}}R)\|_{*}\lesssim(\lambda^{\prime})^{n}$. By Lemma 4, $\mathcal{U}_{(\lambda^{\prime})^{-n}(f^{n})_{*}(\mathrm{dd^{c}}R)}$ is $(C_{6}(\lambda^{\prime})^{-n},\tilde{\kappa}^{n})$-Hölder continuous. By Proposition 2.1, we have
$|\langle d_{q}^{-n}(f^{n})^{*}(S),R\rangle|\lesssim Ad_{q}^{-n}(\lambda^{\prime})^{n}(1+\tilde{\kappa}^{-n}\log^{+}(C_{6}(\lambda^{\prime})^{-n}))\lesssim(\lambda^{\prime\prime}\tilde{\kappa}^{-1}d_{q}^{-1})^{n}.$
@@
We have the following convergence towards the Green currents for all currents in $\mathcal{D}_{q}$. Compared with Proposition 3.1, the convergence is no longer SP-uniform. This is due to the fact that $\|R\|_{\mathcal{C}^{2}}$ is not uniformly bounded for all smooth (kq,kq)(k-q,k-q)-form $R$ with $\|\mathrm{dd^{c}}R\|_{*}\leq 1$.
Corollary 4.6. Let $f$ be as in Proposition 4 and $S$ be a current in $\mathcal{D}_{q}$. If the sequence $n_{i}^{1-m}d_{q}^{-n_{i}}(f^{n_{i}})^{*}\{S\}$ converges to some class $c$, then $n_{i}^{1-m}d_{q}^{-n_{i}}(f^{n_{i}})^{*}(S)$ converges to $T_{c}$.
@@
In the following, we assume $f$ is a holomorphic correspondence on $X$ with simple action on cohomology and $f^{-1}$ is also holomorphic. Suppose $d:=d_{q}(f)$ is the main dynamical degree of $f$. Since we have $(f^{-1})^{*}=f_{*}$ and by duality, we get $d_{s}(f^{-1})=d_{k-s}(f)$ for all $0\leq s\leq k$. Therefore, $f^{-1}$ also has simple action on cohomology with the same main dynamical degree. The limits $T^{+}=\lim_{n\to\infty}d^{-n}(f^{n})^{*}(\omega^{q})$ and $T^{-}=\lim_{n\to\infty}d^{-n}(f^{n})_{*}(\omega^{k-q})$ exist in the sense of currents. We call them the main dynamical Green current. They satisfy the relations $f^{*}(T^{+})=dT^{+}$ and $f_{*}(T^{-})=dT^{-}$. We denote by $c^{+}:=\{T^{+}\}\in H^{q,q}(X,\mathbb{R})$ and $c^{-}:=\{T^{-}\}\in H^{k-q,k-q}(X,\mathbb{R})$. By [19, Lemma 5.4], the intersection number $c^{+}\smile c^{-}$ is strictly positive. Moreover, multiplying $\omega$ by a constant we can assume
$\langle T^{+},T^{-}\rangle=c^{+}\smile\{\omega^{k-q}\}=\{\omega^{q}\}\smile c^{-}=1.$
We also need the following result. See [19, Lemma 5.2].
Lemma 4.7. Let $L^{+}$ denote the line spanned by $c^{+}$. Then exists a hyperplane $K^{+}$ of $H^{q,q}(X,\mathbb{R})$, invariant by $f^{*}$, such that $H^{q,q}(X,\mathbb{R})=L^{+}\oplus K^{+}$ and the spectral radius of $d^{-1}f^{*}$ on $K^{+}$, denoted by $r^{+}$, is strictly less than 1.
In this case, $L^{+}$ is exactly the (strictly) dominant eigenspace defined in the previous section. For any $c\in K^{+}$, we have
$c\smile c^{-}=c\smile d^{-n}(f^{n})_{*}(c^{-})=d^{-n}(f^{n})^{*}(c)\smile c^{-}$
for any $n\in\mathbb{N}$. By letting $n$ goes to infinity, we deduce that $c\smile c^{-}=0$.
Note that the correspondences in the family considered in Corollary 5.2 all have simple action on cohomology. The following gives Part (1) of Theorem 1.
Corollary 4.8. Let $f$ be a holomorphic correspondence on $X$ with $q$-small adjoint multiplicity and simple action on cohomology. Let $S$ be a current in $\mathcal{D}_{q}$ with $\langle T^{-},S\rangle=1$. Then for any smooth test (kq,kq)(k-q,k-q)-form $R$ with $\|R\|_{\mathcal{C}^{2}}\leq 1$, we have
$|\langle d^{-n}(f^{n})^{*}(S)-T^{+},R\rangle|\leq C_{8}\lambda_{1}^{n}$
for some constants $C_{8}>0$ depending only on $S$ and $f$, and $\lambda_{1}=\max\{\lambda_{0},r_{+}\}$.
@@
Since $\{S\}\smile c^{-}=\langle T^{-},S\rangle=1$, we can write $\{S\}=c^{+}+c_{0}$ for some $c_{0}\in K^{+}$ by previous discussion. Therefore, we can find a smooth closed $(q,q)$-form $\beta$ and $S^{\prime\prime}\in\mathcal{D}_{q}^{0}$ such that $\{\beta\}\in K^{+}$ and $S=T^{+}+\beta+S^{\prime\prime};$. It suffices to prove $d^{-n}(f^{n})^{*}(\beta+S^{\prime\prime})$ converges to 0 exponentially fast. The result follows from $\|d^{-n}(f^{n})^{*}(\beta)\|\lesssim r_{+}^{n}$ and applying Proposition 4 to $S^{\prime\prime}$. ∎
@@
Definition 4.11. Let $f$ be a holomorphic correspondence on $X$ such that $f^{-1}$ is also a holomorphic correspondence. We say that $f$ has small multiplicity if the holomorphic correspondence $F=(f,f^{-1})$ has $k$-small adjoint multiplicity.